A project costs $1,400,000, has a four year life, and no re-sale value at the end...

  1. A project costs $1,400,000, has a four year life, and no re-sale value at the end of the project. Depreciation is straight-line to zero. The opportunity cost of capital is 9% and the tax rate is 34%. Sales are projected at 110 units per year. The price per unit is expected to be $30,000, variable costs are projected at $20,500 per unit, and fixed costs will be $515,000 per year. Cash flows from operations are received annually beginning in one year.

    1. (a) What is the NPV?

    2. (b) Conduct a sensitivity analysis with respect to selling price projections (only)

      assuming a pessimistic case and an optimistic case corresponding to plus/minus 10% of the base-case. Assume this project is part of an otherwise profitable corporation that pays corporate tax.

    3. (c) What are the accounting and economic break-even level of sales using expected values?
Answer #1

Intially explained all the data in the question.

Calculating Annual Cash Flows

Discounting the annual cash flows to the present value for calculation of NPV.

Calculating NPV. Answer is $118771.

Calculating NPV of pessimistic and optimistic change in selling price by 10% on base case price side by side in a tabular format.

Calculating Breakeven Sales by accounting perspective where answer is 91 units.

And in calculation of economic breakeven we do back calculation by taking NPV as $0. Answer is 68 units.

Page No. Date Solution: Project Cost Cat time to = $ 1400000 Present Value of cash outflows = $1400 000 Re-sale Value = o Pro

Total Variable Cost (VC) = 20500 * 110 = $ 2255000

Pact . E DE Total fixed cost- $515,000 Cash flow Table Particulars . Amount ($) Sales 3300 000 des Variable cost 2255000 ContPresent Value inflows = Annel, cash inflows & Puff. 4,68,800 X 3.2397 -$15,12,771 Coppor) a Net Present Vanow. Prof inflows-(6) Sensitivity Analysis Particular Pessimistic Optimuste 36000 +104 30000 -10°/. splunit 27500 33000 - 27000 x 110 33ooox 11

% change in NPV are -733.2% and 594.09%

So, by this we can is highly sensitive per unit as with selling price there the NPV. say that the NPV to the selling price 10

(Page No. Date: fixed lost will also incluede depreciation y to tal fivee cast: 515000 + 350oro 2 865000 Break even Sulen 2 8Economici Economic break even es acheived when the NPV of the project is O, NPV = 0 = Prof inflows - Prof outflows Prol inlow

flows Lees Annual cash Depreciation PAT 432139 350000 82139 PBT = PAT/1-2 = 82139 (1-0.34) Add fixed Cost 124453 3515000 - To